OpenStudy (anonymous):

Which is a solution to the equation y = 2x – 5? A. (5, –5) B. (3, 11) C. (1, –3) D. (2, 1)

OpenStudy (anonymous):

What one do you think?

OpenStudy (afrodiddle):

I just passed the lesson you are on! :)

OpenStudy (afrodiddle):

Remember Y= mx + b

OpenStudy (anonymous):

i think a

OpenStudy (afrodiddle):

Do you remember how to graph that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Just plug in your (x,y) into the equation y = 2x-5. For instance, I will do A (5,-5) y = 2x-5 -5=2(5)-5 -5=10-5 -5 = 5 =false so that is not a solution. -5 does not equal 5 so that is not an answer.

OpenStudy (anonymous):

Try B and tell me if it is a solution.

OpenStudy (anonymous):

\[ y = 2x-5\]

OpenStudy (afrodiddle):

Nixy Can I explain this to him?

OpenStudy (anonymous):

Sure :-)

OpenStudy (anonymous):

what makes u think im a him

OpenStudy (afrodiddle):

My bad :P What gender are you?

OpenStudy (anonymous):

i am a man but what if i wasnt

OpenStudy (afrodiddle):

Lol, I just figured because I have never heard a girl say "For days"

OpenStudy (anonymous):

really

OpenStudy (afrodiddle):

Yup.

OpenStudy (anonymous):

I think we are getting side tracked

OpenStudy (afrodiddle):

Any how, have you attended the live lesson for this?

OpenStudy (afrodiddle):

The live lesson helped me a lot. What the moderator went over is pretty much the formula Y=Mx + b

OpenStudy (afrodiddle):

M = y - intercept B = the slope

OpenStudy (anonymous):

there was no live lesson

OpenStudy (afrodiddle):

There was a live lesson yesterday at 7 pm 5.01

OpenStudy (anonymous):

on what

OpenStudy (afrodiddle):

module 5 lesson 1

OpenStudy (anonymous):

The answer will be C. Just plug in your values and be done with it.

OpenStudy (afrodiddle):

Nixy :/

OpenStudy (anonymous):

what site are u from

OpenStudy (afrodiddle):

Your not from Flvs?

OpenStudy (anonymous):

Have fun!

OpenStudy (afrodiddle):

Nixy if you have a skype, I can teach you how to graph, it will make class a lot easier.

OpenStudy (afrodiddle):

BLAH, not Nixy, I meant pizza

OpenStudy (anonymous):

im from k12

OpenStudy (anonymous):

i have skype

OpenStudy (afrodiddle):

My name is GamingRobinHood

OpenStudy (anonymous):

1 sec im gettig in

OpenStudy (anonymous):

i tried to call u